6. The average speed of a nitrogen molecule in air is about

6.70 1 102 m/s, and its mass is 4.68 1 10#26 kg. (a) If it

takes 3.00 1 10#13 s for a nitrogen molecule to hit a wall

and rebound with the same speed but moving in the op-

posite direction, what is the average acceleration of the

molecule during this time interval? (b) What average

force does the molecule exert on the wall?


1
Expert's answer
2021-05-05T16:05:20-0400

Given,

Velocity"(v_1)=6.70\times 10^{2} m/s"

After the hitting the wall, it rebounds with the same speed in opposite direction,

"v_2=-6.7\times 10^{2}m/s"


Mass "(m)=4.68\times 10^{-26}kg"


Time "(t)=3.0\times 10^{-13}s"


"a=\frac{dv}{dt}=\frac{v_2-v_1}{t}"


"=\frac{-6.7\times 10^2-6.7\times 10^2}{3.0\times 10^{-13}} m/s^2"


"=-4.47\times10^{15} m/s^2"

Average force "(F)= ma"


"F=4.68\times 10^{-26}\times 4.47\times 10^{15}N"

"=20.91\times 10^{-11}N"